# Piotr Berman Meiram Murzabulatov

### v1

@@ 3.2 Testing Border-Connectedness
    Let $\mathcal{S}$ be the set of squares of all levels $i\in[0..\log\frac{1}{{\epsilon}})$. By independence and Lemma 3.3, the overall probability that the subroutine accepts in all iterations is at most
    $\prod_{i\in[0..{\log\frac{1}{{\epsilon}}})}\left(1-\frac{\alpha\cdot k_{i}}{2n^{2}}\cdot\sum_{s\in\mathcal{S}_{i}}elc(s)\right)^{2^{i+1}}$ $\leq\prod_{i\in[0..{\log\frac{1}{{\epsilon}}})}\exp\left(-\frac{\alpha}{2}\cdot\frac{4}{{\epsilon}n^{2}2^{i}}\cdot\sum_{s\in\mathcal{S}_{i}}elc(s)\cdot 2^{i+1}\right)$
    $=\exp\left(-\frac{4\alpha}{{\epsilon}n^{2}}\cdot\sum_{s\in\mathcal{S}}els(s)\right)$ $\leq\exp\left(-\frac{4\alpha}{{\epsilon}n^{2}}\cdot\frac{{\epsilon}n^{2}}{2}\right)=e^{-2\alpha}.$
    Therefore, the probability that Algorithm 1 detects at least one witness is at least $1-e^{-2\alpha}$. Since at least an ${\epsilon}$ fraction of pixels in $M$ are black and every square of every level contains at most $\frac{16}{{\epsilon}^{2}}\leq\frac{{\epsilon}n^{2}}{4}$ pixels (recall that $n\geq 8{\epsilon}^{-3/2}$ in the premise of Theorem 1.1), the probability that Algorithm 1 detects a black pixel outside of that witness in Step 1 is at least $1-(1-\frac{3{\epsilon}}{4})^{\frac{4}{{\epsilon}}}>1-e^{-3}$. Thus, for both values of $\alpha$, the probability that Algorithm 1 rejects $M$ is at least
    $(1-e^{-3})(1-e^{-2\alpha})\geq 2/3,$
@@
    We prove that Algorithm 1 has query complexity $O(\frac{1}{{\epsilon}^{2}})$ if it uses Exhaustive-Square-Tester as a subroutine. Algorithm 1 samples $2^{i+1}$ squares of level $i\in[0..\log\frac{1}{{\epsilon}})$ and, for each sampled square, it calls Exhaustive-Square-Tester which makes $(\frac{4}{{\epsilon}}\cdot 2^{-i}-1)(\frac{4}{{\epsilon}}\cdot 2^{-i}-1)<\frac{16}{{\epsilon}^{2}2^{2i}}$ queries in each sampled square of level $i$. Thus, the query complexity of Algorithm 1 is
    $\sum\nolimits_{i=0}^{\log\frac{1}{{\epsilon}}}2^{i+1}\cdot\frac{16}{{\epsilon}^{2}2^{2i}}<\sum\nolimits_{i=0}^{\log\frac{1}{{\epsilon}}}\frac{32}{{\epsilon}^{2}2^{i}}=O\left(\frac{1}{{\epsilon}^{2}}\right).$
    When Algorithm 1 uses Diagonal-Square-Tester, it queries at most $\frac{2k_{i}^{2}}{m_{i}}$ diagonal lattice pixels inside each square of level $i$ (in Step 2 of the subroutine). After that, in Step 2 of the subroutine, it selects $\frac{1}{2}{k_{i}m_{i}}$ pixels and a number $x\in[k_{i}^{2}]$ from the specified distribution and then makes at most $4x$ queries for each selected pixel. Observe that $\operatorname*{\mathbb{E}}(x)=O(\log k_{i})$. Thus, the expected number of queries inside a square of level $i$ is at most $\frac{2k_{i}^{2}}{m_{i}}+\frac{1}{2}k_{i}m_{i}\cdot 4\cdot O(\log k_{i})=O(k_{i}^{3/2}\sqrt{\log k_{i}})$. The expected total number of queries is $\sum\nolimits_{i=0}^{\log(1/{\epsilon})}O(k_{i}^{3/2}\sqrt{\log k_{i}})\cdot 2^{i+1}=O({\epsilon}^{-3/2}\sqrt{\log\frac{1}{{\epsilon}}})$.
@@
    The time complexity of Step 1a and Step 2 of Algorithm 1 is $O(\frac{1}{{\epsilon}})$. Therefore, the total time complexity of Algorithm 1 is $O(\frac{1}{{\epsilon}})$+time complexity of Step 1b. In Step 1b, Algorithm 1 uses either Exhaustive-Square-Tester or Diagonal-Square-Tester. Both of them perform a breadth first search within each sampled square. Breadth first search is linear in the sum of the number of edges and the number of nodes of the graph. Every pixel of a sampled square has at most $4$ neighboring pixels. Thus, the number of edges in the image graph of every sampled square is linear in the number of pixels inside it and the time complexity of Step 1b is linear in the number of all queried pixels, i.e., $O(\frac{1}{{\epsilon}^{2}})$ for Exhaustive-Square-Tester and $O({\epsilon}^{-3/2}\sqrt{\log(1/{\epsilon})})$ for Diagonal-Square-Tester. This completes the proof of the theorem.
@@ 4 Lower Bound for Testing Connectedness
    In this section, we give a lower bound on the query complexity of testing connectedness, proving Theorem 1.2. We use the standard set up of constructing a distribution $\mathcal{N}$ on ${\epsilon}$-far inputs such that every deterministic nonadaptive algorithm that makes $q\leq\frac{c}{{\epsilon}}\log\frac{1}{{\epsilon}}$ queries (for some constant $c$) has probability of error greater than 1/3. By Yao’s Principle [22], it is sufficient to prove Theorem 1.2.
    4.1 Construction of $\mathcal{N}$
    Figure 4.1: Our construction of $\mathcal{N}$: an interesting window together with a column of black pixels immediately to the right of it. All other pixels in the constructed image are white.
    The construction is parameterized by $n$ and the proximity parameter ${\epsilon}$. It gives a distribution supported on $(n+1)\times(n+1)$ images that are ${\epsilon}$-far from connectedness. We assume that ${\epsilon}$ is sufficiently small and that $n>\frac{1}{16}(\frac{1}{{\epsilon}})^{5/8}.$
    We also assume that $n$ is a power of $2$ and $1/{\epsilon}$ is an even power of $2$ and that both of them are sufficiently large, so that all indices in our construction are integer. (See also Footnote 2 for the discussion of integrality issues.) Our construction starts by selecting a level. The indices of the levels range from the low index $\ell=\frac{1}{8}\log\frac{1}{{\epsilon}}$ to the high index $h=\frac{1}{4}\log\frac{1}{{\epsilon}}$. For all $i\in\{\ell,\ell+1,\dots,h\}$, define $a_{i}=2^{i}$ and $n_{i}=16\cdot\sqrt{\epsilon}\cdot n\cdot a_{i}=\sqrt{\epsilon}\cdot n\cdot 2^{i+4}$. First, we pick a uniformly random integer $i\in\{\ell,\ell+1,\dots,h\}$. Consider the $n\times n$ image resulting from removing the last row and the last column of the $(n+1)\times(n+1)$ image. We partition this $n\times n$ image into $(\frac{n}{n_{i}})^{2}$ squares with side length $n_{i}$ called windows of level $i$; that is, each window of level $i$ is an $n_{i}\times n_{i}$ subimage. We pick one of the windows of level $i$ uniformly at random and call it an interesting window. We make the $n_{i}$ pixels immediately to the right of the interesting window black, representing a vertical black line segment, and all other pixels outside of the interesting window white.
@@ 4.1 Construction of $\mathcal{N}$
    We fill the interesting window with a checkerboard pattern of squares of size $a_{i}\times a_{i}$ pixels. Inside each white checkerboard square that is not in the first column, number the rows starting from 0 and make every odd row, excluding the last, fully black, except for one randomly selected pixel for each row. The resulting $\frac{a_{i}}{2}-1$ black lines, each containing one white pixel, are called bridges. The white pixels on the bridges are called disconnecting pixels. We refer to each checkerboard square with the bridges as a bridge square.
    The intuition behind the construction is the following. Each black checkerboard square is in its own connected component. However, to “catch” this connected component as a witness of disconnectedness, a tester would have to query all the disconnecting pixels in the bridge squares to the left and/or to the right of the black square. Since the positions of the disconnecting pixels are random, it would have to query $\Omega(a_{i}^{2})$ pixels in at least one relevant bridge square. Since the windows are selected at random, the algorithm would have to do it for many windows of each level. The key feature of our construction is that interesting windows of different levels are either disjoint or contained in one another, so potential witnesses for one window can’t significantly help with another.
@@ 4.2 Analysis of the Construction
    There are $\frac{(n_{i}/a_{i})^{2}}{2}=128\epsilon n^{2}$ black regions, each one in a separate connected component except for the last $\frac{n_{i}/a_{i}}{2}=8\sqrt{\epsilon}\cdot n$ which are connected by the vertical line. Thus, the image graph has at least $128\cdot\epsilon n^{2}-8\sqrt{\epsilon}\cdot n$ connected components. Changing one pixel from white to black corresponds to adding a node of degree at most 4 to the image graph. This decreases the number of connected components by at most 3. Removing a pixel decreases the number of connected components by at most 1. Consequently, overall, we need to change at least $\frac{1}{3}{(128\epsilon n^{2}-8\sqrt{\epsilon}\cdot n)}$ pixels. This is at least $\epsilon n^{2}$ for sufficiently large $n$; in particular, it holds for $n>\frac{1}{16}(\frac{1}{{\epsilon}})^{5/8}$, which was our assumption in the beginning of Section 4.1. ∎
@@
    We define an event $E$ that must happen in order for the algorithm ${\mathcal{A}}$ to succeed in finding a witness of disconnectedness in an image distributed according to $\mathcal{N}$. Then we show that the probability of $E$ is too small. To define $E$, we define a special group of pixels.
@@
    Since the tester ${\mathcal{A}}$ has 1-sided error, it can reject only if it finds a violation of connectedness. In particular, for an image distributed according to $\mathcal{N}$, in order to succeed, it must find a revealing set for the interesting window of the image.
@@
    An important feature of our construction is that the largest bridge square is smaller than the smallest window. Indeed, the side length of the largest bridge square is $a_{h}=(\frac{1}{{\epsilon}})^{1/4}$, whereas the side length of the smallest window is $n_{\ell}=16\sqrt{{\epsilon}}\cdot(\frac{1}{{\epsilon}})^{1/8}n=16{\epsilon}^{3/8}n$. Thus, $a_{h}<n_{\ell}$ as long as $n>\frac{1}{16}(\frac{1}{{\epsilon}})^{5/8}$, as we assumed in the beginning of Section 4.1. The consequence of this feature and the fact that both $a_{i}$’s and $n_{i}$’s are powers of 2 is that each bridge square of level $i$ is contained in one window of level $j$ for all levels $i$ and $j.$
    A (potential) bridge square of level $i\in\{\ell,\ell+1,\dots,h\}$ is covered if ${\bf Q}$ contains at least ${a_{i}^{2}}/8$ pixels from that square. A window of level $i$ is good if it contains a covered bridge square of level $j\geq i$; otherwise, it is bad. For each good window $w$, we pick a covered bridge square of the highest level contained in $w$ and call it the covered bridge square associated with $w$. All windows of the same level are associated with different covered bridge squares, because each bridge square is contained in exactly one window of a given level.
@@
    Next, we analyze event $G$. For every level $i\in\{\ell,\ell+1,\dots,h\},$ let $g_{i}$ be the number of good windows of level $i$ associated with covered bridge squares of level $i$ and let $t_{i}$ be the total number of good windows of level $i$. Then $g_{h}=t_{h}$ and $g_{i}=t_{i}-t_{i+1}$ for all $i\in[\ell,h-1].$ Observe that, for each level $i\in[\ell,h]$, the covered bridge squares associated with the good windows of level $i$ are distinct. By definition, each of them contributes at least ${a_{i}^{2}}/8$ towards ${\bf Q}$. Therefore, the number of all pixels in ${\bf Q}$ satisfies
    $q$ $\geq\sum_{i=\ell}^{h}\frac{a^{2}_{i}g_{i}}{8}=\frac{1}{8}\left(\sum_{i=\ell}^{h-1}a^{2}_{i}(t_{i}-t_{i+1})+a_{h}^{2}t_{h}\right)=\frac{1}{8}\left(\sum_{i=\ell}^{h-1}4^{i}(t_{i}-t_{i+1})+4^{h}t_{h}\right)$
    $\geq\frac{3}{32}\sum_{i=\ell}^{h}(4^{i}t_{i}).$ (5)
    Recall that $q\leq\frac{c}{{\epsilon}}\log\frac{1}{{\epsilon}}$, that the total number of levels is $h-\ell+1=\Theta(\log\frac{1}{{\epsilon}})$, and that the number of all windows for each level $i$ is $(\frac{n}{n_{i}})^{2}=\Theta(\frac{1}{4^{i}{\epsilon}}).$ Thus,

### v2

@@ 4 Repeat $\frac{1}{2}k_{i}m_{i}$ times:
Let $\mathcal{S}$ be the set of squares of all levels $i\in[0..\log\frac{1}{{\epsilon}})$. Recall that $k_{i}=\frac{4}{{\epsilon}}\cdot 2^{-i}-1$ and $\frac{1}{{\epsilon}}$ is a power of 2. Then, $\log\frac{1}{{\epsilon}}$ is an integer and $i\leq\log\frac{1}{{\epsilon}}-1$. Consequently, $k_{i}\geq\frac{3.5}{{\epsilon}}\cdot 2^{-i}$. By independence, the fact that $1-x\leq e^{-x}$ for all $x$, and Lemma 3.3, the overall probability that the subroutine accepts in all iterations is at most
$\prod_{i\in[0..{\log\frac{1}{{\epsilon}}})}\left(1-\frac{\alpha\cdot k_{i}}{2n^{2}}\cdot\sum_{s\in\mathcal{S}_{i}}elc(s)\right)^{2^{i+1}}$ $\leq\prod_{i\in[0..{\log\frac{1}{{\epsilon}}})}\exp\left(-\frac{\alpha}{2}\cdot\frac{3.5}{{\epsilon}n^{2}2^{i}}\cdot\sum_{s\in\mathcal{S}_{i}}elc(s)\cdot 2^{i+1}\right)$
$=\exp\left(-\frac{3.5\alpha}{{\epsilon}n^{2}}\cdot\sum_{s\in\mathcal{S}}els(s)\right)$ $\leq\exp\left(-\frac{3.5\alpha}{{\epsilon}n^{2}}\cdot\frac{{\epsilon}n^{2}}{2}\right)=e^{-1.75\alpha}.$
Therefore, the probability that Algorithm 1 detects at least one witness is at least $1-e^{-1.75\alpha}$. Since at least an ${\epsilon}$ fraction of pixels in $M$ are black and every square of every level contains at most $\frac{16}{{\epsilon}^{2}}\leq\frac{{\epsilon}n^{2}}{4}$ pixels (recall that $n\geq 8{\epsilon}^{-3/2}$ in the premise of Theorem 1.1), the probability that Algorithm 1 detects a black pixel outside of that witness in Step 1 is at least $1-(1-\frac{3{\epsilon}}{4})^{\frac{8}{{\epsilon}}}>1-e^{-6}$. Thus, for both values of $\alpha$, the probability that Algorithm 1 rejects $M$ is at least
$(1-e^{-6})(1-e^{-1.75\alpha})\geq 2/3,$
@@
Step 1 of Algorithm 1 makes $O(\frac{1}{{\epsilon}})$ queries. We first prove that Algorithm 1 has query complexity $O(\frac{1}{{\epsilon}^{2}})$ if it uses Exhaustive-Square-Tester as a subroutine. In Steps 11, Algorithm 1 samples $2^{i+1}$ squares of each level $i\in[0..\log\frac{1}{{\epsilon}})$ and, for each sampled square, it calls Exhaustive-Square-Tester which makes $(\frac{4}{{\epsilon}}\cdot 2^{-i}-1)(\frac{4}{{\epsilon}}\cdot 2^{-i}-1)<\frac{16}{{\epsilon}^{2}2^{2i}}$ queries in each sampled square of level $i$. Thus, the query complexity of Steps 11 is
$\sum\nolimits_{i=0}^{\log\frac{1}{{\epsilon}}}2^{i+1}\cdot\frac{16}{{\epsilon}^{2}2^{2i}}<\sum\nolimits_{i=0}^{\log\frac{1}{{\epsilon}}}\frac{32}{{\epsilon}^{2}2^{i}}=O\left(\frac{1}{{\epsilon}^{2}}\right),$
and the total query complexity of Algorithm 1 is also $O(\frac{1}{{\epsilon}^{2}}).$
When Algorithm 1 uses Diagonal-Square-Tester, it queries at most $\frac{2k_{i}^{2}}{m_{i}}$ diagonal lattice pixels inside each square of level $i$ (in Step 2 of the subroutine). After that, in Step 2 of the subroutine, it selects $\frac{1}{2}{k_{i}m_{i}}$ pixels and a number $x\in[k_{i}^{2}]$ from the specified distribution and then makes at most $4x$ queries for each selected pixel. Observe that $\operatorname*{\mathbb{E}}[x]=O(\log k_{i})$. Thus, the expected number of queries inside a square of level $i$ is at most $\frac{2k_{i}^{2}}{m_{i}}+\frac{1}{2}k_{i}m_{i}\cdot 4\cdot O(\log k_{i})=O(k_{i}^{3/2}\sqrt{\log k_{i}})$. The expected total number of queries is $\sum\nolimits_{i=0}^{\log(1/{\epsilon})}O(k_{i}^{3/2}\sqrt{\log k_{i}})\cdot 2^{i+1}=O({\epsilon}^{-3/2}\sqrt{\log\frac{1}{{\epsilon}}})$.
@@
To analyze the running time, we assume that individual queries and basic arithmetic operations can be performed in constant time. The time complexity of Step 1 of Algorithm 1 is $O(\frac{1}{{\epsilon}})$. Therefore, the total time complexity of Algorithm 1 is $O(\frac{1}{{\epsilon}})$ plus the time complexity of Steps 11. In Step 1b, Algorithm 1 uses either Exhaustive-Square-Tester or Diagonal-Square-Tester.
Exhaustive-Square-Tester can be implemented by running a BFS on the image graph for the sampled square to find a proof of disconnectedness. BFS takes time that is linear in the sum of the number of edges and the number of nodes of the graph. Every pixel of a sampled square has at most $4$ neighboring pixels. Thus, the number of edges in the image graph of every sampled square is linear in the number of pixels inside it and the time complexity of Steps 11 is linear in the number of all queried pixels, i.e., $O(\frac{1}{{\epsilon}^{2}})$ when Exhaustive-Square-Tester is used.
Next, we argue that Diagonal-Square-Tester takes time linear in the number of pixels it queries. The number of diamonds it considers is at most linear in the number of queried pixels. Step 2, where Diagonal-Square-Tester partitions the diamonds into sets $A$ and $B$, can be implemented by running a BFS on the graph where the nodes correspond to the diamonds, and two diamonds are adjacent if there is a black pixel on their shared border. This BFS takes linear time because the graph has constant degree. In Step 2, we check for every queried black pixel what type of diamond it is in. This check can be performed in $O(1)$ time. If the pixel is in a diamond from set $B$, this pixel is used as part of a BFS in Step 4c on the image graph. As we argued before, a BFS on the image graph takes time linear in the number of pixels queried by the BFS. Thus, the time complexity of Steps 11 of Algorithm 1 is linear in the number of all queried pixels, i.e., $O({\epsilon}^{-3/2}\sqrt{\log(1/{\epsilon})})$ when Diagonal-Square-Tester is used. This completes the proof of Theorem 1.1.
@@ 4 Lower Bound for Testing Connectedness
In this section, we give a lower bound on the query complexity of testing connectedness, proving Theorem 1.2. We use the standard setup of constructing a distribution $\mathcal{N}$ on ${\epsilon}$-far inputs such that every deterministic nonadaptive algorithm that makes $q\leq\frac{c}{{\epsilon}}\log\frac{1}{{\epsilon}}$ queries (for some constant $c$) has probability of error greater than 1/3. Then Theorem 1.2 follows by Yao’s Principle [24].
4.1 Construction of the Hard Distribution
Figure 4.1: Our construction of distribution $\mathcal{N}$: an interesting window together with a column of black pixels immediately to the right of it. All other pixels in the constructed image are white.
The construction is parameterized by a positive integer $n$ and the proximity parameter ${\epsilon}\in(0,1)$. It gives a distribution supported on $(n+1)\times(n+1)$ images that are ${\epsilon}$-far from connectedness. We assume that $n>\frac{1}{16}(\frac{1}{{\epsilon}})^{5/8}.$
We also assume that $n$ is a power of $2$ and $1/{\epsilon}$ is an even power of $2$ and that both of them are sufficiently large, so that all indices in our construction are integers. (See also Footnote 2 for the discussion of integrality issues.) Our construction starts by selecting a level. The indices of the levels range from the low index $\ell=\frac{1}{8}\log\frac{1}{{\epsilon}}$ to the high index $h=\frac{1}{4}\log\frac{1}{{\epsilon}}$. For all indices $i\in\{\ell,\ell+1,\dots,h\}$, define $a_{i}=2^{i}$ and $n_{i}=16\cdot\sqrt{\epsilon}\cdot n\cdot a_{i}=\sqrt{\epsilon}\cdot n\cdot 2^{i+4}$. First, we pick a uniformly random integer $i\in\{\ell,\ell+1,\dots,h\}$. Consider the $n\times n$ image resulting from removing the last row and the last column of the $(n+1)\times(n+1)$ image. We partition this $n\times n$ image into $(\frac{n}{n_{i}})^{2}$ squares with side length $n_{i}$ called windows of level $i$; that is, each window of level $i$ is an $n_{i}\times n_{i}$ subimage. We pick one of the windows of level $i$ uniformly at random and call it an interesting window. We make the $n_{i}$ pixels immediately to the right of the interesting window black, representing a vertical black line segment, and all other pixels outside of the interesting window white.
@@ 4.1 Construction of the Hard Distribution
We fill the interesting window with a checkerboard pattern consisting of $a_{i}\times a_{i}$ squares. Inside each white checkerboard square that is not in the first column, number the rows starting from 0 and make every odd row, excluding the last, fully black, except for one randomly selected pixel for each row. The resulting $\frac{a_{i}}{2}-1$ black lines, each containing one white pixel, are called bridges. The white pixels on the bridges are called disconnecting pixels. We refer to each checkerboard square with the bridges as a bridge square.
The intuition behind the construction is the following. Each black checkerboard square is in its own connected component. However, to “catch” this connected component as a witness of disconnectedness, a tester would have to query all the disconnecting pixels in the bridge squares to the left and/or to the right of the black square. Since the positions of the disconnecting pixels are random, it would have to query $\Omega(a_{i}^{2})$ pixels in at least one relevant bridge square. To catch a witness of disconnectedness for a specific level, the algorithm would have to do it for many windows of that level because the interesting window is positioned randomly. The key feature of our construction is that interesting windows of different levels are either disjoint or contained in one another, so potential witnesses for one window cannot significantly help with another. More precisely, we prove that a tester must query $\Omega(\frac{1}{{\epsilon}})$ pixels per level. Since there are $\Theta(\log\frac{1}{{\epsilon}})$ levels, this leads to the $\Omega(\frac{1}{{\epsilon}}\log{\frac{1}{{\epsilon}}})$ final query complexity.
Note that it is essential that the lower bound is for nonadaptive testers, since an adaptive tester could use $O(\frac{1}{{\epsilon}})$ queries to identify the level and the position of the interesting window, followed by $O(\frac{1}{{\epsilon}})$ queries to check one of the bridges inside the interesting window.
@@ 4.2 Analysis of the Construction
For every $i\in\{\ell,\ldots,h\}$, there are $\frac{(n_{i}/a_{i})^{2}}{2}=128\epsilon n^{2}$ black regions. Each of them is in a separate connected component, except for the last $\frac{n_{i}/a_{i}}{2}=8\sqrt{\epsilon}\cdot n$, which are connected by the vertical line. Thus, the image graph has at least $128\cdot\epsilon n^{2}-8\sqrt{\epsilon}\cdot n$ connected components. Changing one pixel from white to black corresponds to adding a node of degree at most 4 to the image graph. This decreases the number of connected components by at most 3. Removing a pixel decreases the number of connected components by at most 1. Consequently, overall, we need to change at least $\frac{1}{3}{(128\epsilon n^{2}-8\sqrt{\epsilon}\cdot n)}$ pixels. This is at least $\epsilon n^{2}$ for sufficiently large $n$; in particular, it holds for $n>\frac{1}{16}(\frac{1}{{\epsilon}})^{5/8}$, which was our assumption in the beginning of Section 4.1. ∎
@@
For each level $i\in[\ell,h]$, consider the partition of the image into squares with side length $a_{i}$ and call each such square a cell of level $i$. Note that when image $M$ is drawn from $\mathcal{N}$, the bridge squares of level $i$ are selected from cells of level $i$. An important feature of our construction is that the largest cell is smaller than the smallest window. Indeed, the side length of the largest cell is $a_{h}=(\frac{1}{{\epsilon}})^{1/4}$, whereas the side length of the smallest window is $n_{\ell}=16\sqrt{{\epsilon}}\cdot(\frac{1}{{\epsilon}})^{1/8}n=16{\epsilon}^{3/8}n$. Thus, $a_{h}<n_{\ell}$, as long as $n>\frac{1}{16}(\frac{1}{{\epsilon}})^{5/8}$, as we assumed in the beginning of Section 4.1. The consequence of this feature and the fact that both $a_{i}$’s and $n_{i}$’s are powers of 2 is that each cell of level $i$ is contained in one window of level $j$ for all levels $i,j\in\{\ell,\ell+1,\dots,h\}.$
Recall that set $\bf Q$ of queries is fixed, and that bridge squares are chosen randomly when the image $M$ is drawn from $\mathcal{N}$. Intuitively, in order to certify that $M$ is disconnected, the tester has to query many points in some bridge square. To accomplish this, $\bf Q$ might contain many queries from some cells. However, when $\bf Q$ contains a small number of queries (specifically, at most $\frac{c}{{\epsilon}}\log\frac{1}{{\epsilon}}$ queries), only a small number of cells can contain many queries. We classify windows as good or bad based on whether they contain a cell with sufficiently many queries in $\bf Q$. We give up on good windows in the sense that we analyze the failure probability of the algorithm only for the case when a bad window is chosen as an interesting window. We show that the probability of failure for that case is already too high.
Definition 4.1 (Covered cells, good windows).
A cell of level $i\in\{\ell,\ell+1,\dots,h\}$ is covered if ${\bf Q}$ contains at least ${a_{i}^{2}}/8$ pixels from that cell. A window of level $i$ is good if it contains a covered cell of level at least $i$; otherwise, it is bad. [^3] We defined a good window of level $i$ with respect to covered cells of level $i$ or above, as opposed to just level $i$. Observe that if a cell of level $i+1$ is covered, then at least one of the four cells of level $i$ contained in the considered cell is also covered. Consequently, if a window is good, it contains a covered cell of its own level. While our choice of “level $i$ or above” in the definition does not change its substance compared to just “level $i$”, it highlights that larger cells can make a window good. This is important in the description and analysis of the association procedure in the next part of the proof.
The difficulty with counting queries needed for testing images from $\mathcal{N}$ is that some queries can help with covering different cells. To resolve this issue, we define maximal cells and give a procedure that associates many good windows with maximal cells. We call a cell maximal if it is covered and not contained in another covered cell. Now we associate some of the good windows with unique maximal cells. To do this, we go through levels from highest to lowest, and for each level $i\in[\ell,h]$, we consider good windows in an arbitrary order. If for a considered window $w$, all of the maximal cells of level at least $i$ inside $w$ are not associated with any previously considered window, we associate $w$ with an (arbitrary) one of those maximal cells. This association allows us to avoid overcounting queries needed when considering different levels. All windows of the same level are associated with different covered cells, because each cell is contained in exactly one window of a given level.
Claim 4.3 (Guarantees of the Association Procedure).
For every level $i\in\{\ell,\ell+1,\dots,h\},$ let $G_{i}$ be the set of good windows of level $i$ associated with maximal cells, let $T_{i}$ be the set of all good windows of level $i$, and define $g_{i}=|G_{i}|$, and $t_{i}=|T_{i}|.$ Then the following hold:
  1. 1.
    $q\geq\sum_{i=\ell}^{h}\frac{a^{2}_{i}g_{i}}{8}$, where $q=|{\bf Q}|.$
  2. 2.
    $g_{h}=t_{h}$ and $g_{i}\geq t_{i}-t_{i+1}$ for all $i\in[\ell,h-1].$
Proof.
We prove the two items separately.
  1. 1.
    For all distinct levels $i,j\in[\ell,h]$, the sets $G_{i}$ and $G_{j}$ are disjoint because they contain windows of different levels. All windows in $\bigcup_{i\in[\ell,h]}G_{i}$ are associated with different maximal cells by construction of associations. Moreover, for all levels $i\in[\ell,h]$, each window in $G_{i}$ is associated with a maximal cell of level at least $i$, that is, a covered cell of that level. By definition, such a cell contributes at least $a^{2}_{i}/8$ queries to $\bf Q.$
  2. 2.
    To prove $g_{h}=t_{h}$, note that each maximal cell of level $h$ is contained in exactly one window of $G_{h}$. Thus, all windows in $T_{h}$ are associated with maximal cells, implying $G_{h}=T_{h}$, and consequently $g_{h}=t_{h}$.
    For the next part of the proof, we use the following terminology. For every window $w$ of a level $i\in[\ell,h-1]$, the parent of $w$, denoted $p_{w}$, is the window of level $i+1$ that contains $w$. Window $w$ is called a child of $p_{w}$. In our construction, each window of level $i\in[\ell+1,h]$ has 4 children.
    Now we fix a level $i\in[\ell,h-1]$ and argue that $t_{i}\leq g_{i}+t_{i+1}$. The sets $G_{i}$ and $T_{i+1}$ are disjoint because they contain windows of different levels. Therefore, $|G_{i}\cup T_{i+1}|=g_{i}+t_{i+1}.$ To prove $t_{i}\leq g_{i}+t_{i+1}$, we establish an injection from $T_{i}$ to $G_{i}\cup T_{i+1}$. Each window of $T_{i}$ that is also in $G_{i}$ is mapped to itself, and each window $w$ in $T_{i}\setminus G_{i}$ is mapped to $p_{w}$, i.e., the parent of $w$.
    Next, we assume that $w\in T_{i}\setminus G_{i}$ and argue that $p_{w}\in T_{i+1}$. Since $w$ is a good window of level $i$, it contains a covered cell of level at least $i$. Recall that the largest cell is smaller than the smallest window, and that a window cannot contain only a part of a cell (because the side lengths of windows and cells in our partitions of the image are powers of two). As a result, $w$ also contains a maximal cell. Since $w$ is in $T_{i}$, but not in $G_{i}$, one of the maximal cells from $w$ is associated with some window $w^{\prime}$ that was considered before $w$ in the construction of associations. Call this maximal cell $c_{w}$. Since each cell belongs to exactly one window of level $i$, window $w^{\prime}$ must be of level at least $i+1.$ Therefore, $c_{w}$ is a covered cell of level at least $i+1$. Since cell $c_{w}$ belongs to window $w$, it also belongs to window $p_{w}$. Consequently, $p_{w}$ is good because it is a window of level $i+1$ and contains a covered cell of level at least $i+1$. So, $p_{w}\in T_{i+1}.$
    Finally, we show that our mapping from $T_{i}$ to $G_{i}\cup T_{i+1}$ is injective. Consider windows $w_{1},w_{2}\in T_{i}$. Suppose for the sake of contradiction that they are mapped to the same window $w$ in $G_{i}\cup T_{i+1}$. Since $w_{1}$ and $w_{2}$ are windows of the same level, and we map each of them either to itself or its parent, both $w_{1}$ and $w_{2}$ must be children of $w$. Since $w_{1}$ and $w_{2}$ were mapped to their parent, they are not associated with any maximal cells. It means that both of them contain maximal cells associated with windows of higher levels. Let $c_{1}$ and $c_{2}$ be such maximal cells in $w_{1}$ and in $w_{2}$, respectively. Suppose w.l.o.g. that $c_{1}$ was associated first. Let $w^{\prime}$ be the window associated with $c_{2}$. The window $w^{\prime}$ must be of level at least $i+1$, so it contains both $c_{1}$ and $c_{2}$. But the procedure that constructed associations could associate $w^{\prime}$ with $c_{2}$ only if all maximal cells of the currently considered level inside $w^{\prime}$ are not associated with any of the previously considered windows. Since $c_{1}$ was already associated, we get a contradiction.∎
@@
Next, we define an event $E$ that must happen in order for the algorithm ${\mathcal{A}}$ to succeed in finding a witness of disconnectedness in an image distributed according to $\mathcal{N}$. Then we show that the probability of $E$ is too small.
@@
Since the tester ${\mathcal{A}}$ has 1-sided error, it can reject only if it finds a violation of connectedness. In particular, to succeed on an image distributed according to $\mathcal{N}$, algorithm ${\mathcal{A}}$ must find a revealing set for the interesting window of the image.
@@
Next, we analyze the event $G$.
We begin by giving a lower bound on $q$, the number of queries in $\bf Q$ in terms of the number of good windows of different levels. We first apply Items 1 and 2 of 4.3 and then rearrange the terms of the resulting summation:
$q$ $\geq\sum_{i=\ell}^{h}\frac{a^{2}_{i}g_{i}}{8}\geq\frac{1}{8}\left(\sum_{i=\ell}^{h-1}a^{2}_{i}(t_{i}-t_{i+1})+a_{h}^{2}t_{h}\right)$
$=\frac{1}{8}\left(\sum_{i=\ell}^{h-1}4^{i}(t_{i}-t_{i+1})+4^{h}t_{h}\right)=\frac{1}{8}\left(\sum_{i=\ell}^{h}4^{i}t_{i}-\sum_{i=\ell+1}^{h}4^{i-1}t_{i}\right)$
$\geq\frac{3}{32}\sum_{i=\ell}^{h}4^{i}t_{i}.$ (5)
Recall that $q\leq\frac{c}{{\epsilon}}\log\frac{1}{{\epsilon}}$, that the total number of levels is $h-\ell+1=\Theta(\log\frac{1}{{\epsilon}})$, and that the number of all windows for each level $i$ is $(\frac{n}{n_{i}})^{2}=\Theta(\frac{1}{4^{i}{\epsilon}}).$ When $M$ is drawn from $\mathcal{N}$, first a level $i$ is chosen uniformly at random, i.e., each level is selected with probability $\frac{1}{h-\ell+1}$, and then an interesting window is chosen uniformly at random among the $(\frac{n}{n_{i}})^{2}$ available windows of level $i$. Since $t_{i}$ windows of level $i$ are good, the probability that a good window is chosen as an interesting window is